OpenStudy (anonymous):

(Can someone work out this problem so I know how to do this type of problem?) Find the average rate of change of f(x) =-x^3 + 1: a) from 0 to 2 b) from 1 to 3 c) from -1 to 1

OpenStudy (ash2326):

@HeyMolli I'll show you how to find the average rate of change for a You do it for b and c then :)

OpenStudy (ash2326):

Average Rate of f(x) from x=a to b is defined as \[Rate=\frac{f(b)-f(a)}{b-a}\] we have \[f(x)=-x^3+1\] a) 0 to 2 Let's find f(2), put x=2 in f(x) \[f(2)=-2^3+1=-8+1=-7\] \[f(0)=-0+1=1\] so \[Rate=\frac{f(2)-f(0)}{2-0}\ \ \ \ \ \ \ \ \ \ \ =\frac{-7-1}{2-0} \ \ \ \ \ \ =-4\] so average rate of change is -4 Do you understand this?

OpenStudy (anonymous):

Yes! Thank you!

OpenStudy (ash2326):

Could you try to find the rate for the other two options?

OpenStudy (anonymous):

okay, i think b is -9.6, and c is -1.

OpenStudy (ash2326):

check b again, c is correct

OpenStudy (anonymous):

Okay, well, if f(3) is -26 and f(1) is zero, then it's -26/(3-1), which is -26/2, so -13. Maybe?

OpenStudy (ash2326):

yeah :) good

OpenStudy (anonymous):

awesome, thanks so much!